Qd=200-1/2×4910+1/4×5900-5×90+1/5×55000
1
Expert's answer
2018-12-12T10:00:11-0500

Qd = 200 - 1/2×4910 + 1/4×5900 - 5×90 + 1/5×55000 = 200 - 2455 + 1475 - 450 + 11000 = 9770.